# On A Scale From 1 to 5: Quantifying Hallucination in Faithfulness Evaluation

### v2

@@ Abstract
Hallucination has been a popular topic in natural language generation (NLG). In real-world applications, unfaithful content can result in bad data quality or loss of trust from end users. Thus, it is crucial to fact-check before adopting NLG for production usage, which can be expensive if done manually. In this paper, we investigate automated faithfulness evaluation in guided NLG. We developed a rubrics template and use large language models (LLMs) to score the generation into quantifiable scales. We compared popular LLMs as well as the widely adopted natural language inference (NLI) models in scoring quality and sensitivity. In addition, we developed methods to generation synthetic unfaithful data, as well as a heuristics to quantify the percentage of hallucination. Our results on 4 travel-domain industry dataset show that GPT-4 (Achiam et al., 2023) can provide accurate judgement and explanation on whether a source and a generation are factually consistent. Furthermore, we found that tuning NLI models on synthetic data can improve performance. Lastly, we present insights on latency and cost for deploying such system.
@@
Xiaonan Jing, Srinivas Billa,  and Danny Godbout Expedia Group Inc. {xijing, nbilla, dgodbout}@expediagroup.com
@@ 1 Introduction
The recent advancement of large language models (LLM) has made it easier than ever to build NLP applications. With LLM’s powerful natural language generation (NLG) ability, one can easily leverage prompt engineering to generate texts. However, pretrained transformer decoders are prone to hallucinations. Previously it was shown that the performance of LLM’s decoding strategies (e.g. beam search) have a dependency on the type of the generation task (Wiher et al., 2022). As a result, LLMs can sometimes generate completed "made-up" sequences, in open-ended generation like story telling and close-ended generation like summarization (Maynez et al., 2020). Many recent methods have been proposed to mitigate hallucination, including retrieval-augmented generation (Lewis et al., 2020), chain-of-thoughts (Wei et al., 2022), tree-of-thoughts (Yao et al., 2024), self-consistency (Wang et al., 2023b), self-reflection (Ji et al., 2023b), controlled decoding (Mudgal et al., 2023), instruction tuning (Liu et al., 2023a). Nevertheless, the current state-of-the-art does not guarantee that hallucination can be 100% prevented. To apply NLG in industry applications, oftentimes having a "faithful enough" model is not sufficient to productionize an LLM-driven experience. With the aim of adding another layer of security to assure the accuracy of LLM generated content, we investigate an automated approach to effectively score degrees of faithfulness in guided text generation.
@@
Figure 1 illustrates a simple example of guided generation scenarios. Given guidelines (the prompt template) and grounding data (the source facts), an LLM model can hallucinate and generate either intrinsic hallucinating content in which the facts contradict that of the source or extrinsic hallucinating content in which new facts are added that cannot be verified from the source (Ji et al., 2023a). It should be noted that extrinsic hallucination does not necessarily contain incorrect information with respect to world knowledge. It is possible that the LLM interpreted the additional facts based on its internal knowledge. On the other hand, faithfulness is considered an antonym to "hallucination", which is defined as staying factually consistent with the provided source (Ji et al., 2023a). A higher faithfulness score thus indicates lower hallucinations. We define faithfulness of a generated text with respect to its source as the following:
@@
The hypothesis is factually consistent with the reference if all factual elements in the hypothesis can be traced back to the reference. A faithfulness score measures: 1) to what extent the hypothesis can be verified by the reference; and 2) to what extent the LLM hallucinates when generating the hypothesis from the reference.
We investigated two approaches, namely using LLMs with a grading rubrics and natural language inference (NLI) models, to derive a faithfulness score for a given reference and hypothesis pair. We experimented with 5 LLMs and 3 variants of NLI models on 4 travel-domain dataset. We studied the score variation by comparing various types of hypothesis: gold hypothesis, sentence-level gold hypothesis, intrinsic hypothesis, and extrinsic hypothesis. In addition, we explored the model sensitivity in picking up different levels of hallucination. Our main contributions are: First, we showed that LLMs can accurately distinguish faithfulness apart from hallucination, and that reasoning during grading help LLMs gain sensitivity towards hallucination. Second, we presented methods to generate synthetic hallucination by type and demonstrated that tuning with synthetic data can improve NLI model’s performance. Lastly, we illustrated the progression of scores based on the percentage of hallucinating content.
@@ 2 Related Works
NLI for NLG evaluation NLI based models has been used widely in factual consistency evaluation. Derived originally from the text entailment (Dagan et al., 2005), given a premise, a model should classify the premise to be an "entailment", "neutral", or "contradiction" to a hypothesis (Bowman et al., 2015). Laban et al. (2022) proposed SummaC (Laban et al., 2022) which demonstrated that a fact check with NLI-based models on summarization performed better when using sentence-level granularity in both reference and hypothesis. Honovich et al. (2022) introduced TRUE benchmark for factual consistency evaluation, which showed that the NLI-based binary classifier could retain state-of-the-art performance over summarization, dialogue, paraphrasing, and fact checking tasks. Falsesum (Utama et al., 2022) is a data generation pipeline that utilized OpenIE and a fine-tuned T5-base (Raffel et al., 2020) model to create synthetic factually inconsistent data for improving NLI-based classifier. The "neutral" and "contradiction" labels were together treated as "non-entailment".
LLM for NLG evaluation. Recently, many studies have also proven that an LLM could be used as a judge for evaluating NLP tasks (Zheng et al., 2023). Wang et al. (2023a) showed that ChatGPT can perform as well as human judges in NLG evaluation, in which a broader generation quality was graded between the reference and the hypothesis. Liu et al. (2023b) proposed G-Eval which used a chain-of-thought framework to evaluate NLG by granular categories. The faithfulness component was drafted as a binary classification for factual inconsistency identification. Chiang and Lee (2023) expanded on G-Eval and demonstrated that grading with explanation can improve the LLMs evaluation accuracy in summarization. Although the factual consistency was scored on a 1-5 scale, LLM’s internal knowledge was used as criteria. Chang et al. (2024) showed that LLM can detect hallucinations as a binary classification problem in RAG systems in a zero-shot setting, and that training on synthetic data can further improve the detection accuracy. Gekhman et al. (2023) presented TrueTeacher which improved both LLM- and NLI-based factual consistency model accuracy by training on LLM-generated synthetic datasets.
Previous works on NLI models focus on classification and lack of explainability, while LLM-based evaluation is still emerging which has not arrived formalization. As far as we know, few studies have tried to quantify the degree of faithfulness. Furthermore, little documentation about how NLI models compare to LLM evaluation is available for domain specific real-world use cases. Finally, comparison in terms of what percentage of hallucinations affect the final scores has not been explored extensively.
@@ 3 System
We evaluated faithfulness scoring models on 4 travel-domain industry datasets. It should be noted that our datasets are taken from production and only contains gold data. We discuss means to create synthetic hallucination dataset in Section 3.1. Statistics of the dataset can be found in Table 1.
@@ 3.1 Dataset
ConvoAS is an abstractive summarization dataset. Our source data consists of customer support transcripts between 2 participants, a traveler and an agent. GPT-4o (OpenAI, 2024b) was used to generate a summary for the entire transcript. Information in the summary such as date, price, and location needs to be traceable from the source.
ConvoTS is a topic-specific summarization dataset. Similar to ConvoAS, this dataset also contains customer support transcripts, except 3 topic-specific summaries were generated for each transcript. It should be noted that many redundant turns exist because the chat happened between a traveler and a virtual agent through a guided service. GPT-4o (OpenAI, 2024b) was used for this use case.
ReviewTS is a topic-specific summarization dataset from reviews. Each reference consists of multiple review snippets related to a topic of interest (e.g. pool). A fine-tuned Mistral 7B (Jiang et al., 2023) model was used to extract and summarize the snippets. The summary is restricted to less than 100 characters. Because of the concise nature of this data, full-length and sentence-level hypothesis variants contain large overlaps. However, we kept the same setup for consistency.
@@
Synthetic Hallucination Data. Due to insufficient hallucination examples, we drafted guidelines for GPT-4o to generate a synthetic counterpart from the gold dataset to study model performance. Since the hallucination could take the form of either intrinsic, for contradictory facts, or extrinsic for extra knowledge, we mimicked both scenarios. For the former, we prompted GPT-4o to modify a piece of fact in the original sentence; and for the latter, we instructed the model to incorporate a piece of new world knowledge while the original sentence remains unchanged. Each dataset thus resulted in 4 varying hypothesis versions, namely gold full-length, gold sentence-level, intrinsic sentence-level, and extrinsic sentence-level. Furthermore, we simulated a numeric percentage at the sentence level to measure how much effect the portion of unfaithful content has on the evaluation scores. We took a ConvoAS subset with a 5-sentence hypothesis (51 samples). For every gold hypothesis, we incrementally swapped out 1 to 5 random sentences by its counterpart. The percentage of sentences that contain unfaithful information can thus be treated to mimic the "hallucination percentage" with respect to the overall hypothesis. It should be noted that it is inherently difficult to justify "the number of facts" contained in downstream tasks like abstractive summarization because the granularity of the content is subjective to the use case and the boundary between different "pieces of facts" can be fuzzy. While the proposed heuristic approach is far from perfection, it is sufficient to serve our goal of learning the scoring trend and model sensitivity.
@@ 3.2 Entailment-based Evaluation
We adopted Vectara’s open-source HHEMv1.0111https://github.com/vectara/hallucination-leaderboard/tree/hhem-1.0-final model (Bao et al., 2023). HHEM is a 184.4 million parameter lightweight Deberta model (He et al., 2021) fine-tuned on various NLI dataset, including TRUE (Honovich et al., 2022) and SummaC (Laban et al., 2022). Given a pair of text (reference, hypothesis), HHEM scores the level of entailment between the pair to a value in the range [0, 1], with 1 being entailment and 0 being contradictory.
One limitation is that HHEM model has a max input size of 512 tokens, which is prone to long text pairs. By default, truncation is applied to overflowing tokens starting at the end of the texts. In this case, the hypothesis will be truncated first, which will lead to information loss and worse case, totally wrong evaluation. To minimize information loss for maximized accuracy, we implemented a sliding window segmentation to the reference. Overlapping tokens were taken into consideration to include the previous context. The size of the current segment is computed dynamically based on tokens in the hypothesis as shown below:
@@
The hypothesis is then evaluated against every reference segment and a final score for each pair is max aggregated across all segments. The reason behind segmenting the reference rather than the hypothesis is that the reference is often longer and contains redundancy. It is less sensible to segment the hypothesis, in which the information is typically already compressed, as it will result in more data loss. For the size of the overlapping tokens, we intuitively set it to 32 tokens which corresponds to around 25 words. This length is slightly longer than the length of an average sentence, which is said to be around 15-20 words.
@@ 3.3 Rubrics-based LLM Evaluation
Following the accuracy metric defined in HELM Liang et al. (2022), we crafted a rubrics-based approach to use LLM as a judge to score the faithfulness of the generated contents.
@@
We defined the grading rubric to follow a discrete range from 1 to 5, with 5 being "highly faithful" that all information in the hypothesis can be verified in the reference, and 1 being "highly unfaithful" which most or all of the content in the hypothesis cannot be validated from the reference. Additionally, as prior research has shown that self-reflection by including an explanation (Renze and Guven, 2024) can effectively improve LLM performance, we also instructed the model to provide detailed reasoning along with the numeric score to further improve the grading precision. More details of the prompt can be found in Appendix A.1
@@ 4 Result & Discussion
In this section, we discuss faithfulness scoring on the reference and hypothesis pairs for two sets of experiments. The score ranges from "1-highly unfaithful", "2-very faithful", "3-slightly unfaithful", and "4-very unfaithful", to "5-highly faithful". It should be noted that HHEM models return a continuous probabilistic value in the range [0, 1]. It was later scaled to [1, 5] for ease of interpretation.
@@
Our first experiment aims to evaluate the models’ scoring quality on gold and hallucinating contents. The experiments (Figure 2) outline model comparison across all datasets over the gold hypothesis, gold hypothesis on a sentence level, intrinsic hypothesis on a sentence level, and extrinsic hypothesis on a sentence level. On the other hand, the second set of experiments aims to study the score progression when the percentage of hallucinating content varies in the hypothesis. The comparison (Figure 3) involves scoring on both intrinsic and extrinsic hypotheses with 0% to 100% hallucinating contents in 20% steps. Along with the score progression, we also tested the significance of "reasoning" to LLM’s scoring capability. Models tested are as follows: for rubrics-based scoring with LLMs: GPT-4 (Achiam et al., 2023), GPT-4o-mini (OpenAI, 2024a), Llama3.1-8B Meta (2024), Llama3.1-70B Meta (2024), and Claude3 Haiku Anthropic (2024a); for entailment-based approach HHEM (Bao et al., 2023), HHEM-intrinsic, and HHEM-extrinsic.

### v3

@@ Abstract
Hallucination has been a popular topic in natural language generation (NLG). In real-world applications, unfaithful content can result in poor data quality or loss of trust from end users. Thus, it is crucial to fact-check before adopting NLG for production usage, which can be expensive if done manually. In this paper, we investigate automated faithfulness evaluation in guided NLG. We developed a rubric template and used large language models (LLMs) to score the generation on quantifiable scales. We compared popular LLMs as well as widely adopted natural language inference (NLI) models in scoring quality and sensitivity. In addition, we developed methods for the generation of synthetic unfaithful data, as well as heuristics to quantify the percentage of hallucination. Our results on 4 travel-domain industry dataset show that GPT-4 (Achiam et al., 2023) can provide accurate judgement and explanation of whether a source and a generation are factually consistent. Furthermore, we found that tuning NLI models on synthetic data can improve performance. Lastly, we present insights on the latency and cost of deploying such a system.
@@
Xiaonan Jing, Srinivas Billa,  and Danny Godbout Expedia Group Inc. $1$ Expedia Group Wy W Seattle, WA 98119, USA {xijing, nbilla, dgodbout}@expediagroup.com
@@ 1 Introduction
The recent advancement of large language models (LLM) has made it easier than ever to build NLP applications. With LLM’s powerful natural language generation (NLG) ability, one can easily leverage prompt engineering to generate texts. However, pretrained transformer decoders are prone to hallucinations. Previously, it was shown that the performance of LLM’s decoding strategies (e.g., beam search) has a dependency on the type of the generation task (Wiher et al., 2022). As a result, LLMs can sometimes generate completed "made-up" sequences, in open-ended generation like story telling, and close-ended generation like summarization (Maynez et al., 2020). Many recent methods have been proposed to mitigate hallucination, including retrieval-augmented generation (Lewis et al., 2020), chain-of-thoughts (Wei et al., 2022), tree-of-thoughts (Yao et al., 2024), self-consistency (Wang et al., 2023b), self-reflection (Ji et al., 2023b), controlled decoding (Mudgal et al., 2023), instruction tuning (Liu et al., 2023a). Nevertheless, the current state-of-the-art does not guarantee that hallucination can be 100% prevented. To apply NLG in industry applications, oftentimes having a "faithful enough" model is not sufficient to productionize an LLM-driven experience. With the aim of adding another layer of security to ensure the accuracy of LLM generated content, we investigate an automated approach to effectively score degrees of faithfulness in guided text generation.
@@
Figure 1 illustrates a simple example of guided generation scenarios. Given guidelines (the prompt template) and grounding data (source facts), an LLM model can hallucinate and generate either intrinsic hallucinating content in which the facts contradict those of the source or extrinsic hallucinating content in which new facts are added that cannot be verified from the source (Ji et al., 2023a). It should be noted that extrinsic hallucination does not necessarily contain incorrect information with respect to world knowledge. It is possible that the LLM interpreted the additional facts based on its internal knowledge. On the other hand, faithfulness is considered an antonym for "hallucination", which is defined as staying factually consistent with the provided source (Ji et al., 2023a). A higher faithfulness score thus indicates lower hallucinations. We define the faithfulness of a generated text with respect to its source as the following:
@@
The hypothesis is factually consistent with the reference if all the factual elements of the hypothesis can be traced back to the reference. A faithfulness score measures: 1) to what extent the hypothesis can be verified by the reference; and 2) to what extent the LLM hallucinates when generating the hypothesis from the reference.
We investigated two approaches, namely using LLMs with a grading rubric and natural language inference (NLI) models, to derive a faithfulness score for a given reference and hypothesis pair. We experimented with 5 LLMs and 3 variants of NLI models on 4 travel-domain dataset. We studied the score variation by comparing various types of hypothesis: gold hypothesis, sentence-level gold hypothesis, intrinsic hypothesis, and extrinsic hypothesis. In addition, we explored the model sensitivity in picking up different levels of hallucination. Our main contributions are as follows. First, we showed that LLMs can accurately distinguish faithfulness apart from hallucination, and that reasoning during grading helps LLMs gain sensitivity towards hallucination. Second, we presented methods to generate synthetic hallucinations by type and demonstrated that tuning with synthetic data can improve the performance of the NLI model. Lastly, we illustrated the progression of scores based on the percentage of hallucinating content.
@@ 2 Related Works
NLI for NLG evaluation NLI based models has been used widely in factual consistency evaluation. Derived originally from the text entailment (Dagan et al., 2005), given a premise, a model should classify the premise as an "entailment", "neutral" or "contradiction" to a hypothesis (Bowman et al., 2015). Laban et al. (2022) proposed SummaC (Laban et al., 2022) which demonstrated that a fact check with NLI-based summarization models performed better when using sentence-level granularity in both reference and hypothesis. Honovich et al. (2022) introduced TRUE benchmark for factual consistency evaluation, which showed that the NLI-based binary classifier could retain state-of-the-art performance over summarization, dialogue, paraphrasing, and fact checking tasks. Falsesum (Utama et al., 2022) is a data generation pipeline that utilized OpenIE and a fine-tuned T5-base (Raffel et al., 2020) model to create synthetic factually inconsistent data for improving NLI-based classifier. The "neutral" and "contradiction" labels were together treated as "non-entailment".
LLM for NLG evaluation. Recently, many studies have also proven that an LLM could be used as a judge to evaluate NLP tasks (Zheng et al., 2023). Wang et al. (2023a) showed that ChatGPT can perform as well as human judges in NLG evaluation, in which a broader generation quality was graded between the reference and the hypothesis. Liu et al. (2023b) proposed G-Eval which used a chain-of-thought framework to evaluate NLG by granular categories. The faithfulness component was drafted as a binary classification for the identification of factual inconsistency. Chiang and Lee (2023) expanded on G-Eval and demonstrated that grading with explanation can improve the LLMs evaluation accuracy in summarization. Although factual consistency was scored on a scale of 1-5, LLM’s internal knowledge was used as a criterion. Chang et al. (2024) showed that LLM can detect hallucinations as a binary classification problem in RAG systems in a zero-shot setting and that training on synthetic data can further improve detection accuracy. Gekhman et al. (2023) presented TrueTeacher which improved both LLM- and NLI-based factual consistency model accuracy by training on LLM-generated synthetic datasets.
Previous works on NLI models focus on classification and lack of explainability, while LLM-based evaluation is still emerging, which has not arrived at formalization. As far as we know, few studies have tried to quantify the degree of faithfulness. Furthermore, little documentation about how NLI models compare to LLM evaluation is available for domain-specific real-world use cases. Finally, a comparison in terms of the percentage of hallucinations that affect the final scores has not been explored extensively.
@@ 3 System
We evaluated faithfulness scoring models on 4 travel-domain industry datasets. It should be noted that our datasets are taken from production and only contain gold data. We discuss means to create synthetic hallucination datasets in Section 3.1. The statistics of the dataset can be found in Table 1.
@@ 3.1 Dataset
ConvoAS is an abstractive summarization dataset. Our source data consists of customer support transcripts between 2 participants, a traveler, and an agent. GPT-4o (OpenAI, 2024b) was used to generate a summary of the entire transcript. The information in the summary, such as date, price, and location, should be traceable from the source.
ConvoTS is a topic-specific summarization dataset. Similarly to ConvoAS, this dataset also contains customer support transcripts, except 3 topic-specific summaries were generated for each transcript. It should be noted that many redundant turns exist because the chat happened between a traveler and a virtual agent through a guided service. GPT-4o (OpenAI, 2024b) was used for this use case.
ReviewTS is a topic-specific summarization dataset from reviews. Each reference consists of multiple review snippets related to a topic of interest (e.g., pool). A fine-tuned Mistral 7B (Jiang et al., 2023) model was used to extract and summarize the snippets. The summary is restricted to less than 100 characters. Because of the concise nature of these data, the full-length and sentence-level hypothesis variants contain large overlaps. However, we kept the same setup for consistency.
@@
Synthetic Hallucination Data. Due to insufficient hallucination examples, we drafted guidelines for GPT-4o to generate a synthetic counterpart from the gold dataset to study model performance. Since the hallucination could take the form of either intrinsic, for contradictory facts, or extrinsic, for extra knowledge, we mimicked both scenarios. For the former, we prompted GPT-4o to modify a piece of fact in the original sentence; and for the latter, we instructed the model to incorporate a piece of new world knowledge while the original sentence remains unchanged. Each dataset thus resulted in 4 varying hypothesis versions, namely gold full-length, gold sentence-level, intrinsic sentence-level, and extrinsic sentence-level. Furthermore, we simulated a numeric percentage at the sentence level to measure how much effect the portion of unfaithful content has on the evaluation scores. We took a ConvoAS subset with only 5-sentence hypothesis (51 samples). For every gold hypothesis, we incrementally swapped out 1 to 5 random sentences by its counterpart. The percentage of sentences that contain unfaithful information can thus be treated to mimic the "hallucination percentage" with respect to the overall hypothesis. It should be noted that it is inherently difficult to justify "the number of facts" contained in downstream tasks like abstractive summarization because the granularity of the content is subjective to the use case, and the boundary between different "pieces of facts" can be fuzzy. While the proposed heuristic approach is far from perfect, it is sufficient to serve our goal of learning the scoring trend and model sensitivity.
@@ 3.2 Entailment-based Evaluation
We adopted Vectara’s open source HHEMv1.0111https://github.com/vectara/hallucination-leaderboard/tree/hhem-1.0-final model (Bao et al., 2023). HHEM is a 184.4 million parameter lightweight Deberta model (He et al., 2021) fine-tuned on various NLI datasets, including TRUE (Honovich et al., 2022) and SummaC (Laban et al., 2022). Given a pair of text (reference, hypothesis), HHEM scores the level of entailment between the pair to a value in the range [0, 1], with 1 being entailment and 0 being contradictory.
One limitation is that the HHEM model has a maximum input size of 512 tokens, which is prone to long text pairs. By default, truncation is applied to overflowing tokens starting at the end of the texts. In this case, the hypothesis will be truncated first, which will lead to information loss and worse case to a totally wrong evaluation. To minimize information loss for maximum accuracy, we implemented sliding window segmentation in the reference. Overlapping tokens were taken into consideration to include the previous context. The size of the current segment is computed dynamically based on tokens in the hypothesis as shown below:
@@
The hypothesis is then evaluated against every reference segment, and a final score for each pair is max aggregated across all segments. The reason behind segmenting the reference rather than the hypothesis is that the reference is often longer and contains redundancy. It is less sensible to segment the hypothesis, in which the information is typically already compressed, as it will result in more data loss. For the size of the overlapping tokens, we intuitively set it to 32 tokens which corresponds to around 25 words. This length is slightly longer than the length of an average sentence, which is said to be around 15-20 words.
@@ 3.3 Rubrics-based LLM Evaluation
Following the accuracy metric defined in HELM Liang et al. (2022), we developed a rubric-based approach to use LLM as a judge to score the faithfulness of the generated content.
@@
We defined the grading rubric to follow a discrete range from 1 to 5, with 5 being "highly faithful" that all information in the hypothesis can be verified in the reference and 1 being "highly unfaithful" that most or all of the content in the hypothesis cannot be validated from the reference. Furthermore, since previous research has shown that self-reflection by including an explanation (Renze and Guven, 2024) can effectively improve LLM performance, we also instructed the model to provide detailed reasoning along with the numeric score to further improve the grading precision. More details of the prompt can be found in the Appendix A.1
@@ 4 Result & Discussion
In this section, we discuss the faithfulness scoring on the reference and hypothesis pairs for two sets of experiments. The score ranges from "1-highly unfaithful", "2-very faithful", "3-slightly unfaithful", and "4-very unfaithful" to "5-highly faithful". It should be noted that HHEM models return a continuous probabilistic value in the range [0, 1]. It was later scaled to [1, 5] for ease of interpretation.
@@
Our first experiment aims to evaluate the model scoring quality on gold and hallucinating contents. The experiments (Figure 2) outline model comparison across all datasets on the gold hypothesis, the gold hypothesis on a sentence level, the intrinsic hypothesis on a sentence level and the extrinsic hypothesis on a sentence level. On the other hand, the second set of experiments aims to study the progression of the score when the percentage of hallucinating content varies according to the hypothesis. The comparison (Figure 3) involves scoring both intrinsic and extrinsic hypotheses with 0% to 100% hallucinating contents in 20% steps. Along with the score progression, we also tested the significance of "reasoning" to LLM’s scoring capability. The models tested are as follows: for rubric-based scoring with LLMs: GPT-4 (Achiam et al., 2023), GPT-4o-mini (OpenAI, 2024a), Llama3.1-8B Meta (2024), Llama3.1-70B Meta (2024), and Claude3 Haiku Anthropic (2024a); for the entailment-based approach, HHEM (Bao et al., 2023), HHEM-intrinsic, and HHEM-extrinsic.